Will you please help me solve the following quadratic equation?
m=(m[0])/sqrt(1-(v^2/c^2)) solving for c
This is what I've attempted...
multiply both sides by the LCM sqrt1-(v^2/c^2)to get rid of the fraction
(m)qrt1-(v^2/c^2)=m[0]
then multiply both sides again to get rid of the fraction in the sqrt?
(mc^2)sqrtc^2-v^2=m[0]
then divide both sides by (m)sqrt(c^2-v^2)
c^2=(m[0])/(m)sqrt(c^2-v^2)
square both sides with a final answer of
c=sqrt(m[0])/(m)sqrt(c^2-v^2)
Thank you in advance for your help!!
Brenda

Answer by stanbon(75887)

m=(m[0])/sqrt(1-(v^2/c^2)) solving for c
-----------------
Interchange m and [sqrt(1-(v^2/c^2)) to get:
sqrt(1-(v^2/c^2)) = m[0]/m
-----------------------
Square both sides to get
1-(v^2/c^2) = [m[0]/m]^2
-------------
Then v^2/c^2 = 1-[m(0]/m]^2
--------------
And c^2 == v^2/[1-[m[0]/m]^2]
-------------
Take the sqrt to find "c".
================
Cheers,
Stan H.